what is the equation of the line with a slope of -4/3 and containg the point (6,-2)?
Answer by stanbon(75887)

what is the equation of the line with a slope of -4/3
and containingg the point (6,-2)?
------------------------
Form: y = mx + b
Substitute for x,y, and m ; solve for "b":
-2 = (-4/3)*6 + b
-2 = -8 + b
b = 6
-------
Equation:
y = (-4/3)x + 6
=================
Cheers,
Stan H.
=================